LOG-IN
Displaying report 1-1 of 1.
MIF (ITF Control)
Print this report.
yutaro.enomoto - 16:45, Monday 03 December 2018 (7265)Get code to link to this report
Frequency noise of the laser injected to MIF were evaluated by X arm

Using the data taken yesterday's lock of X arm, the external disturbance of this loop, (which is dominated by frequency fluctuation of main laser at f >~1 Hz), are quickly plotted here along with the estimated frequency noise of IMC transmission.

NOTE:
-- This estimation of IOO frequency noise was done for the Phase-1 configuration, which is different from current one. So these two traces cannot be compared directly.
-- No precise fitting for OLTF has not yet been done. The OLTF of CARM loop was modelled just using the measured unitiy gain frequency of ~2.5kHz (at this moment, IN1 gain was not 31dB, but 26dB. So UGF was not 5kHz, but 2.5kHz).

Images attached to this report
7265_20181203084519_ioofreqnoisecomparisonxarm.png
Comments related to this report
yutaro.enomoto - 22:27, Wednesday 05 December 2018 (7283)

[Miyazaki, Sugimoto, Enomoto]

We measured PD dark noise and it is added to the IOO frequency noise evaluation plot.